### v1

# Explicit modeling of density dependence in spatial capture-recapture models

Abstract
Density dependence occurs at the individual level but is often evaluated at the population level, leading to difficulties or even controversies in detecting such a process. Bayesian individual-based models such as spatial capture-recapture (SCR) models provide opportunities to study density dependence at the individual level, but such an approach remains to be developed and evaluated. In this study, we developed a SCR model that links habitat use to apparent survival and recruitment through density dependent processes at the individual level. Using simulations, we found that the model can properly inform habitat use, but tends to underestimate the effect of density dependence on apparent survival and recruitment. The reason for such underestimations is likely due to the fact that SCR models have difficulties in identifying the locations of unobserved individuals while assuming they are uniformly distributed. How to accurately estimate the locations of unobserved individuals, and thus density dependence, remains a challenging topic in spatial statistics and statistical ecology.
Keywords: population regulation, Bayesian individual-based modeling, landscape heterogeneity, habitat selection, vital rate, demography
1 Introduction
Identifying the effect of density dependence in population regulation is a fundamental question in ecology and often relies on statistical approaches (St. Amant,, 1970; Hanski,, 1990; Turchin and Taylor,, 1992; Hassell,, 1975; Herrando-Pérez et al.,, 2012). The answer to this question has broad implications for evolution and theoretical ecology (Bellows,, 1981; Ratikainen et al.,, 2008; Travis et al.,, 2013; Holt,, 1996) and application values in predicting population viability and extension under habitat changes (Howell et al.,, 2020; Zhao et al.,, 2019; Zabel et al.,, 2006; Sabo et al.,, 2004). While most studies examined density dependence at the population level (Dennis and Taper,, 1994; Brook and Bradshaw,, 2006; Dennis et al.,, 2006; Hanski,, 1990), processes that lead to density dependence, such as competition (Lamb et al.,, 2017; Hixon and Jones,, 2005; Fronhofer et al.,, 2023), often occurs at the individual level. Such discrepancy can lead to difficulties in detecting density dependence and even controversial results and imposes challenges in statistical modeling (Detto et al.,, 2019; Shenk et al.,, 1998).
Individual-based modeling has been used to understand how population-level patterns emerge from individual-level processes (Judson,, 1994; Grimm and Railsback,, 2013). While individual-based modeling has traditionally relies on supervised learning (Grimm and Railsback,, 2013), more recently researchers have emerged it with reinforcement learning under a Bayesian framework (i.e., Bayesian individual-based modeling Hooten et al.,, 2020; Schafer et al.,, 2022). Such a development provides opportunities to gain knowledge about individual-level processes, including density dependence across the individual and population levels in heterogeneous landscapes, with statistical modeling.
Spatial capture-recapture (SCR) was originally developed to estimate population size while assuming a “selection-free” kernel for habitat use (Efford,, 2004; Borchers and Efford,, 2008; Royle and Young,, 2008). These models were then further extended to allow habitat use related to landscape characteristics at local scales (Royle et al.,, 2013, 2018). Dynamic SCR models allow vital rates (apparent survival, recruitment) related to population density, but often at the population scale (Gardner et al.,, 2010; Chandler and Clark,, 2014; Ergon and Gardner,, 2014, but see Milleret et al.,, 2023). Because SCR models utilize individual-based data and explicitly consider the unique location, and thus environment, of each individual, there is a potential to relate vital rates to density dependence at the individual level in these models, along the lines of Bayesian individual-based modeling. However, SCR models that examine the effect of density dependence on vital rates at the individual level remains to be developed and evaluated.
Here, we developed a novel SCR model that estimates habitat use information and links such information to density dependent apparent survival and recruitment rates at the individual level. We evaluated the model with simulations and discussed the pros and cons of our modelling approach and SCR models in general.
2 Methods
2.1 Model description
We developed a dynamic spatial capture-recapture (SCR) models that contains four sub-models: a sub-model that describes habitat use, a sub-model that links habitat use with individual existence through apparent survival and recruitment, a sub-model that describes dispersal movement, and eventually an observation sub-model that link habitat use with frequency of detection data.
The first sub-model describes the amount of time of habitat use by individual animals such that
$\upsilon_{g,i,t}=U\times\frac{\exp(-\kappa d_{g,i,t}+\beta^{[\upsilon]}_{1}x_{i,t,1}+\ldots+\beta^{[\upsilon]}_{n_{covs}}x_{i,t,n_{covs}})}{\sum_{i=1}^{n_{site}}\exp(-\kappa d_{g,i,t}+\beta^{[\upsilon]}_{1}x_{i,t,1}+\ldots+\beta^{[\upsilon]}_{n_{covs}}x_{i,t,n_{covs}})}.$ (1)
Like other SCR models (e.g., Efford,, 2004; Royle et al.,, 2018), we assumed that every individual has a territory with a (time varing) centroid. However, our model is different from previous SCR models in modelling a continuous variable $\upsilon_{g,i,t}$, which represents the true but latent amount of time of individual $g$ using site $i$ in year $t$. More specifically, we assumed that $\upsilon_{g,i,t}$ is a function of the distance between individual $g$’s territory centroid and site $i$, denoted $d_{g,i,t}$, with a decay parameter $\kappa$, as well as $n_{covs}$ potentially time varying environmental covariates $x$’s, with corresponding parameters $\beta$’s representing the direction and strength of habitat selection. To note, because we assume that all possible sites an individual can use are modelled, $\upsilon_{g,i,t}$ add up to $U$, a known quantity for the total amount of time covered by the survey. To satisfy this assumption, the study area needs to be large enough and also relatively isolated so that it contains a distinguishable population; only individuals in this population use the sites in this area. Further, to define these sites, researchers can either rely on prior knowledge about what a site is, or consider every spatial units (e.g., grid cells) in the study area as possible sites and let the model determine if they are used by the animals or not.
The second sub-model describes how the true amount of time of habitat use influences individual existence through density dependent apparent survival and recruitment such that the status of an individual $g$ in year $t$ $z_{g,t}$ (0 for not alive and 1 for alive) follows
$z_{g,t=1}\sim\operatorname{Bernoulli}(\psi)$ (2)
$z_{g,t\geq 2}\sim\operatorname{Bernoulli}\left[z_{g,t-1}\phi_{g,t-1}+(1-e_{g,t-1})\gamma_{g,t-1}\right]$ (3)
$\phi_{g,t}=\operatorname{logit}^{-1}\left[\beta^{[\phi]}_{0}+\beta^{[\phi]}_{density}\times\frac{\sum_{i=1}^{n_{site}}\upsilon_{g,i,t}\left(\sum_{k:k\neq g}\upsilon_{k,i,t}z_{k,t}\right)}{\sum_{i=1}^{n_{site}}\upsilon_{g,i,t}}\right]$ (4)
$\gamma_{g,t}=\operatorname{logit}^{-1}\left[\beta^{[\gamma]}_{0}+\beta^{[\gamma]}_{density}\times\frac{\sum_{i=1}^{n_{site}}\upsilon_{g,i,t}\left(\sum_{k:k\neq g}\upsilon_{k,i,t}z_{k,t}\right)}{\sum_{i=1}^{n_{site}}\upsilon_{g,i,t}}\right]$ (5)
in which $z_{g,t}$ represents the status of individual $g$ in year $t$ such that $z_{g,t}=1$ if individual $g$ exists (i.e., is alive) in year $t$ and $z_{g,t}=0$ otherwise; $\psi$ is the probability of existence in the first year; from the second year, each individual has a probability of $\phi_{g,t-1}$ to survive given it is alive in the previous year (i.e., $z_{g,t-1}=1$), and a probability of $\gamma_{g,t-1}$ to be recruited to the population given it hasn’t existed so far (i.e., $e_{g,t-1}=0$); $e_{g,t}$ is an indicator which equals 1 if individual $g$ has been alive in any of the years up until year $t$, and 0 otherwise; $\phi_{g,t-1}$ and $\gamma_{g,t-1}$ are functions of local density with intercept and slope parameters $\beta^{[\phi]}$’s and $\beta^{[\gamma]}$’s, respectively. The key component of this model, which distinguishes it from other SCR models, is to aggregate habitat use by any individuals other than $g$, $\upsilon_{-g,i,t}$, given they are alive, i.e., $z_{-g,t}=1$, and calculate their weighted mean, using habitat use of individual $g$, $\upsilon_{g,i,t}$, as the weight, across all sites. This approach allows us to represent density dependent processes due to competition such that an individual that more frequently uses habitat sites that are less frequently used by other individuals tends to face a lower level of competition and thus has higher apparent survival and recruitment.
The third sub-model describes dispersal movement as a autoregressive process such that
$\bm{s}_{g,t}\sim\mathcal{N}\left(\bm{s}_{g,t-1},\sigma\bm{I}\right)$ (6)
in which $\bm{s}_{g,t}=(s_{g,t}^{[x]},s_{g,t}^{[y]})$ is the two dimensional vector of individual $g$’s territory centroid in year $t$ with easting $s_{g,t}^{[x]}$ and northing $s_{g,t}^{[y]}$. $\bm{I}$ is two dimensional identity matrix while $\sigma$ is the standard deviation of this Gaussian dispersal movement. Here, we assume that there is no preferred direction of dispersal thus we assume an isotropic Gaussian dispersal kernel. This assumption can be easily extended by allowing correlations between $s_{g,t}^{[x]}$ and $s_{g,t}^{[y]}$.
Eventually, the last sub-model describes how detection data are obtained given the true amount of time of habitat use such that
$y_{g,i,t,j}\sim\operatorname{Poisson}(r_{i,t,j}o_{i,t}z_{g,t}\upsilon_{g,i,t})$ (7)
$r_{i,t,j}=\exp\left(\alpha_{0}+\sum_{k=1}^{n_{\mathrm{pcvs}}}\alpha_{k}w_{i,t,j,k}\right)$
in which discrete data $y_{g,i,t,j}$ represents the number of times for which individual $g$ is captured at site $i$ in year $t$ by the $j$th camera, $r_{i,t,j}$ is the detection rate and is a function of $n_{\mathrm{pcvs}}$ observation covariates $w$’s with intercept and slope parameters $\alpha$’s, and $o_{i,t}$ indicates the proportion of time for which a camera trap is deployed at site $i$ in year $t$. Note that we assumed to have multiple camera traps deployed at each site to allow replicates of the observations (Pollock,, 1982); these camera traps need to function independently, meaning that capturing the animal by one camera trap should not influence the function of other camera traps. Also note that $r_{i,t,j}$ represents average number of detections, rather than a probability, when an activate individual spent unit time at a site that had unit survey effort $o_{i,t}$, and thus is not necessarily upper bounded by 1. Also, because we modelled all possible sites, $o_{i,t}$ allows the situation that camera traps are not necessarily deployed at every site and they can be deployed for different amount of time.
2.2 Simulation design
During the simulations, we simulated a total of 200 individuals, including individuals that actually existed and individuals that did not exist, meaning that we used data augmentation (Royle et al.,, 2007; Lewis et al.,, 2023). We simulated a landscape with 500 sites within a 20-unit wide and 16-unit long area during a 6-year period of time; the landscape was characterized by two spatially and temporally varying covariates, both of which were randomly generated from a standard Normal distribution. We assumed that territory centroids were uniformly distributed in the landscape in the first year. We set $\kappa$ to 2.5 to allow a moderate size of territory. We set $\beta^{[\upsilon]}_{1}$ to 0.6 and $\beta^{[\upsilon]}_{2}$ to -0.3 to allow opposite effects of the two covariates on habitat use. We set $\psi$ to 0.4, meaning that more than half of the individuals did not exist in the first year of the study. We set $\beta^{[\phi]}_{0}$ to 1 and $\beta^{[\gamma]}_{0}$ to -1 to allow a relatively high apparent survival and relatively low recruitment. We set $\beta^{[\phi]}_{density}$ to -0.4 and $\beta^{[\gamma]}_{density}$ to -0.6 to allow negative density dependence on both apparent survival and recruitment. We set $\sigma$ to 2.5 to allow a relatively short distance of dispersal movement. For the observation processes, we considered two observation covariates which again were randomly generated from a standard Normal distribution. We set $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$ to 0.7, -0.6, 0.3, respectively. For cameras in the middle part (16-unit wide and 12-unit long) of the study area, $o_{i,t}$ was randomly generated from a uniform distribution between 0.8 and 1; $o_{i,t}$ was set to 0 to represent the situation that cameras were not deployed outside the middle part of the study area.
We considered two models in our simulations. We first considered a simpler model which only allows density dependence on recruitment but not apparent survival. This is equivalent to have $\phi_{g,t}=\operatorname{logit}^{-1}(\beta^{[\phi]}_{0})$ in equation 2. This model also does not allow dispersal movement between years, which mean we have $\bm{s}_{g,t}=\bm{s}_{g,}$ for any $t$. We then considered a more complex model that allows density dependence on both recruitment and apparent survival, and dispersal movement between years, as described above.
2.3 Model evaluation
We developed custom MCMC algorith in R (R Core Team,, 2024). We repeated the simulations for 20 times with newly generated data each time. We ran 50,000 iterations, including 25,000 burn-in, in 3 chains for each simulation. We pooled all posterior samples from all simulations to make violin plots, which also provided key statistics including median, 50% Credible Interval (CI) and 95% CI. We considered a parameter estimate to be biased if the 50% CI does not cover the true value of the parameter used in the simulation.
2.4 Real data analysis
We applied our proposed method on ten years tiger survey data from India by Karanth et al., (2006). The published version of the dataset by Gardner et al., (2020) contained ten years of survey, however the first two surveys were spaced only half a year. We therefore considered this data to have nine primary period excluding the first snapshot. Survey effort, measured by days in each primary period a camera was active, ranged from 33 to 197 days (detailed in Karanth et al.,, 2006, Table 1). We used these active days to calculate our effort data $o_{i,t}$ dividing days active by 365. We set total number of time covered by survey $U$ to be 1, as measured in year. Within each primary period, there was only one secondary period. There were in total 76 individuals seen thus we augment our data to have 200 potential individuals. The published version of the dataset does not contain any environmental variables and the spatial locations are altered to conserve tiger locations which made it hard to obtain environmental variables through georeferencing, therefore we did not include them in our analysis. Geometric locations were converted to have unit of kilometers instead of original meters. We ran three chains with 50,000 iteractions in each chain.
3 Results
3.1 Simulation
The simulations showed that both the simpler model that only allows density dependence on recruitment but not apparent survival and does not allow dispersal movement between years and the more complex model that allows density dependence on both recruitment and apparent survival and dispersal movement between years can recover parameters related to habitat use ($\kappa$, $\beta^{[\upsilon]}$’s), initial existence ($\psi$) and detection ($\alpha$’s) without evident bias (Figure 1). However, both models underestimated the effects of density dependence on apparent survival ($\beta_{density}^{[\phi]}$) and recruitment ($\beta_{density}^{[\gamma]}$). The simpler model provided unbiased estimates for the mean of apparent survival ($\beta_{0}^{[\phi]}$) and recruitment ($\beta_{0}^{[\gamma]}$) at a logit scale, but the more complex model tended o have slightly biased estimates for these two parameters. The models also underestimated the standard deviation of Gaussian dispersal movement ($\sigma$).
Figure 1: Violin plots showing the combined posterior samples from 20 simulations for the parameters of (M1) a SCR model that only allows density dependence on recruitment but not apparent survival and does not allow dispersal movement between years, and (M2) a SCR model that allows density dependence on both recruitment and apparent survival and dispersal movement.
3.2 Real data
MCMC chains mixed well with Rhat statistics being 1.00 for all parameters of interest. Fig. 2 showed posterior distributions of model parameters. Using our density dependent model, we found a positive density dependence in survival $\phi$ at 90% credible level. This might be a consequence of not having environmental covariates in the model.
Figure 2: Posterior distribution of model parameters estimated using tiger dataset. We found positive density dependence in survival at 90% credible level.
4 Discussion
In this study, we extended SCR models (e.g., Royle et al.,, 2013, 2018) to estimate the true amount of time of habitat use, an information that is often of interest in ecological research and conservation planning (Hull et al.,, 2016; Dawson et al.,, 2013). In our simulations, we used the robust sampling design to facilitate the inference of habitat use. However, the inference of habitat use does not rely on robust sampling design, which is illustrated by our case study.
Figure 3: The true (blue solid point: observed individual, blue broken point: unobserved individual) and estimated territory centroids (yellow cloud: observed individual, red cloud: unobserved individual) in one year of a spatial capture-recapture study. Grey crosses represent the locations of camera traps.
Despite the accurate information of habitat use provided by this SCR model, we found it still difficult to estimate the effect of density dependence on apparent survival and recruitment with this model. So far, density dependence is mostly considered at the population level in SCR models (e.g., Chandler and Clark,, 2014); when density dependence is considered at individual level in SCR models, only the locations of territory centroids, but not the information of habitat use, were considered (Milleret et al.,, 2023). The shortcoming of doing so is that individuals will not influence each other through density dependent processes if their territory centroids do not fall at the same site, despite the fact that their territories can still overlap with each other to induce competition. Therefore, we attempted to develop a model that links habitat use information with density dependence. The difficulties in estimating the effect of density dependence in this model is likely due to the fact that the current generation of SCR models can only identify the locations of territory centroids of observed individuals, but not unobserved individuals, under the assumption of uniformly distributed territory centroids (Figure 3). Because the locations of unobserved individuals are misidentified, the model mis-specifies local density, and thus underestimates the effect of density dependence on vital rates. It is likely for the same reason that the model also provides biased estimates of the movement parameter. Overall, our results are in line with the findings of previous studies that showed the limits of SCR models (Sun and Cole Burton,, 2024), and it remains a challenge to link habitat use with vital rates at individual level in SCR models. Further studies are needed to better specify the locations of individuals, probably through the integration of other data resources with SCR data (e.g., Sun et al.,, 2019).
Our study nonetheless showed the importance of separating true habitat use with observation processes. In our case study, we found an unexpected positive density dependence, likely due to missing environmental covariates because higher density is likely to be an indication of better habitat and in turn higher survival rate. Indeed, including environmental covariates to inform habitat use is a key component of this model. Therefore, we highly recommend future studies to consider habitat heterogeneity in SCR models.