### v2

# Explicit modeling of density dependence in spatial capture-recapture models

Abstract
Density dependence occurs at the individual level and thus is greatly influenced by spatial local heterogeneity in habitat conditions. However, density dependence is often evaluated at the population level, leading to difficulties or even controversies in detecting such a process. Bayesian individual-based models such as spatial capture-recapture (SCR) models provide opportunities to study density dependence at the individual level, but such an approach remains to be developed and evaluated. In this study, we developed a SCR model that links habitat use to apparent survival and recruitment through density dependent processes at the individual level. Using simulations, we found that the model can properly inform habitat use, but tends to underestimate the effect of density dependence on apparent survival and recruitment. The reason for such underestimations is likely due to the difficulties of the current model in identifying the locations of unobserved individuals without using environmental covariates to inform these locations. How to accurately estimate the locations of unobserved individuals, and thus density dependence, remains a challenging topic in spatial statistics and statistical ecology.
Keywords: population regulation, Bayesian individual-based modeling, landscape heterogeneity, habitat selection, vital rate, demography
1 Introduction
Identifying the effect of density dependence in population regulation is a fundamental question in ecology and often relies on statistical approaches (st1970detection; hanski1990density; turchin1992complex; hassell1975density; herrando2012density). The answer to this question has broad implications for evolution and theoretical ecology (bellows1981descriptive; ratikainen2008density; travis2013evolution; holt1996adaptive) and application values in predicting population viability and extension under habitat changes (howell2020informing; zhao2019integrated; zabel2006interplay; sabo2004efficacy). While most studies examined density dependence at the population level (dennis1994density; brook2006strength; dennis2006estimating; hanski1990density), processes that lead to density dependence, such as competition (lamb2017influence; hixon2005competition; fronhofer2023shape), often occur at the individual level and thus is greatly influenced by spatial local heterogeneity in habitat conditions. Such discrepancy can lead to difficulties in detecting density dependence and even controversial results and imposes challenges in statistical modeling (detto2019bias; shenk1998sampling).
Individual-based modeling has been used to understand how population-level patterns emerge from individual-level processes (judson1994rise; grimm2013individual). While individual-based modeling has traditionally relies on supervised learning (grimm2013individual), more recently researchers have emerged it with reinforcement learning under a Bayesian framework (i.e., Bayesian individual-based modeling hooten2020statistical; schafer2022bayesian). Such a development provides opportunities to gain knowledge about individual-level processes, including density dependence across the individual and population levels in heterogeneous landscapes, with statistical modeling.
Spatial capture-recapture (SCR) was originally developed to estimate population size while assuming a “selection-free” kernel for habitat use (efford2004density; borchers2008spatially; royle2008hierarchical; royle2013spatial; efford2025secr). These models were then further extended to allow habitat use to be related to landscape characteristics at local scales (royle2013integrating; royle2018unifying; zhao2024bayesian). Dynamic SCR models allow vital rates (apparent survival, recruitment) to be related to population density, but often at the population scale (gardner2010spatially; chandler2014spatially; ergon2014separating, but see milleret2023estimating). Because SCR models utilize individual-based data and explicitly consider the unique location, and thus environment, of each individual, there is a potential to relate vital rates to density dependence at the individual level in these models, along the lines of Bayesian individual-based modeling. However, SCR models that examine the effect of density dependence on vital rates at the individual level remains to be developed and evaluated.
Here, we developed a novel SCR model that estimates habitat use information and links such information to density dependent apparent survival and recruitment rates at the individual level. We evaluated the model with simulations and a case study, and discussed the pros and cons of our modelling approach and SCR models in general.
2 Methods
2.1 Model description
We developed a dynamic spatial capture-recapture (SCR) model that contains four sub-models: a sub-model that describes habitat use, a sub-model that links habitat use with individual existence through apparent survival and recruitment, a sub-model that describes dispersal movement, and eventually an observation sub-model that link habitat use with frequency of detection data (Figure 1).
Figure 1: Diagram showing the structure of the SCR model introduced in the current study. The model utilizes SCR data $y$, alongside with environmental covariate data $x$, observation covariate data $w$, and survey information $o$, given survey period $U$, to inform habitat use $\upsilon$ with the effect of environment $\beta^{[\upsilon]}$ and the effect of distance $\kappa$ on habitat use, individual existence status $z$ driven by the probabilities of initial existence $\psi$, apparent survival $\phi$ and recruitment $\gamma$ with the effects of habitat use on density dependent apparent survival $\beta^{[\phi]}$ and recruitment $\beta^{[\gamma]}$, the locations of territory centroids $s$ with the standard deviation of Gaussian dispersal $\sigma$, and detection rate $r$ with the effect of observation covariates on detection $\alpha$.
The first sub-model describes the amount of time of habitat use by individual animals such that
$\upsilon_{g,i,t}=U\times\frac{\exp(-\kappa d_{g,i,t}+\beta^{[\upsilon]}_{1}x_{i,t,1}+\ldots+\beta^{[\upsilon]}_{n_{covs}}x_{i,t,n_{covs}})}{\sum_{i=1}^{n_{site}}\exp(-\kappa d_{g,i,t}+\beta^{[\upsilon]}_{1}x_{i,t,1}+\ldots+\beta^{[\upsilon]}_{n_{covs}}x_{i,t,n_{covs}})}.$ (1)
Like other SCR models (e.g., efford2004density; royle2018unifying), we assumed that every individual has a territory with a centroid. However, our model is different from previous SCR models in modelling a continuous variable $\upsilon_{g,i,t}$, which represents the true but latent amount of time of individual $g$ using site $i$ during the primary period (e.g., breeding season) in year $t$, and is assumed to be a function of the distance between individual $g$’s territory centroid and site $i$, denoted $d_{g,i,t}$, with a decay parameter $\kappa$, as well as $n_{covs}$ potentially time varying environmental covariates $x$’s, with corresponding parameters $\beta$’s representing the direction and strength of habitat selection. We also assumed that $\upsilon_{g,i,t}$ add up to $U$, a known quantity for the total amount of time of the primary period covered by the survey, under the assumption of closed population. To satisfy this assumption, the study area needs to be large enough and also relatively isolated so that it contains a distinguishable population, and there should be no changes in the population size due to birth, death, or dispersal movement within each primary period. To satisfy this assumption, researchers also need to model all sites that individuals in this population could possibly use, regardless if the sites are surveyed or not. To define these sites, researchers can either rely on prior knowledge about what a site is, or consider every spatial units (e.g., grid cells) in the study area as possible sites and let the model determine if they are used by the animals or not.
The second sub-model describes how the true amount of time of habitat use influences individual existence through density dependent apparent survival and recruitment such that the status of an individual $g$ in year $t$ $z_{g,t}$ (0 for not alive and 1 for alive) follows
$z_{g,t=1}\sim\operatorname{Bernoulli}(\psi)$ (2)
$z_{g,t\geq 2}\sim\operatorname{Bernoulli}\left[z_{g,t-1}\phi_{g,t-1}+(1-e_{g,t-1})\gamma_{g,t-1}\right]$ (3)
$\phi_{g,t}=\operatorname{logit}^{-1}\left[\beta^{[\phi]}_{0}+\beta^{[\phi]}_{density}\times\frac{\sum_{i=1}^{n_{site}}\upsilon_{g,i,t}\left(\sum_{k:k\neq g}\upsilon_{k,i,t}z_{k,t}\right)}{\sum_{i=1}^{n_{site}}\upsilon_{g,i,t}}\right]$ (4)
$\gamma_{g,t}=\operatorname{logit}^{-1}\left[\beta^{[\gamma]}_{0}+\beta^{[\gamma]}_{density}\times\frac{\sum_{i=1}^{n_{site}}\upsilon_{g,i,t}\left(\sum_{k:k\neq g}\upsilon_{k,i,t}z_{k,t}\right)}{\sum_{i=1}^{n_{site}}\upsilon_{g,i,t}}\right]$ (5)
in which $z_{g,t}$ represents the status of individual $g$ in year $t$ such that $z_{g,t}=1$ if individual $g$ exists (i.e., is alive) in year $t$ and $z_{g,t}=0$ otherwise; $\psi$ is the probability of existence in the first year; from the second year, each individual has a probability of $\phi_{g,t-1}$ to survive given it is alive in the previous year (i.e., $z_{g,t-1}=1$), and a probability of $\gamma_{g,t-1}$ to be recruited to the population given it hasn’t existed so far (i.e., $e_{g,t-1}=0$); $e_{g,t}$ is an indicator which equals 1 if individual $g$ has been alive in any of the years up until year $t$, and 0 otherwise; $\phi_{g,t-1}$ and $\gamma_{g,t-1}$ are functions of local density with intercept and slope parameters $\beta^{[\phi]}$’s and $\beta^{[\gamma]}$’s, respectively. Thus, while we assumed a closed population within each primary period, we allowed changes in the population size between primary periods, which is consistent with many other open-population models (broms2016dynamic; efford2020spatial; howell2020informing; zhao2020sampling; milleret2023estimating). The key component of this model, which distinguishes it from other SCR models, is to aggregate habitat use by any individuals other than $g$, $\upsilon_{-g,i,t}$, given they are alive, i.e., $z_{-g,t}=1$, and calculate their weighted mean, using habitat use of individual $g$, $\upsilon_{g,i,t}$, as the weight, across all sites. This approach allows us to represent density dependent processes due to competition such that an individual that more frequently uses habitat sites that are less frequently used by other individuals tends to face a lower level of competition and thus has higher apparent survival and recruitment.
The third sub-model describes dispersal movement as an autoregressive process such that
$\bm{s}_{g,t}\sim\mathcal{N}\left(\bm{s}_{g,t-1},\sigma\bm{I}\right)$ (6)
in which $\bm{s}_{g,t}=(s_{g,t}^{[x]},s_{g,t}^{[y]})$ is the two dimensional vector of individual $g$’s territory centroid in year $t$ with easting $s_{g,t}^{[x]}$ and northing $s_{g,t}^{[y]}$. Again, we assumed that the territory centroids are constant within each primary period, and only allowed the changes of their locations between primary periods. Also, by having a two dimensional identity matrix $\bm{I}$ with the standard deviation of Gaussian dispersal movement $\sigma$ to represent an isotropic Gaussian dispersal kernel, we assumed that there is no preferred direction of dispersal. This assumption can be easily extended by allowing correlations between $s_{g,t}^{[x]}$ and $s_{g,t}^{[y]}$.
Eventually, the last sub-model describes how detection data are obtained given the true amount of time of habitat use such that
$y_{g,i,t,j}\sim\operatorname{Poisson}(r_{i,t,j}o_{i,t}z_{g,t}\upsilon_{g,i,t})$ (7)
$r_{i,t,j}=\exp\left(\alpha_{0}+\sum_{k=1}^{n_{\mathrm{pcvs}}}\alpha_{k}w_{i,t,j,k}\right)$
in which discrete data $y_{g,i,t,j}$ represents the number of times for which individual $g$ is captured at site $i$ in year $t$ by the $j$th camera, $r_{i,t,j}$ is the detection rate and is a function of $n_{\mathrm{pcvs}}$ observation covariates $w$’s with intercept and slope parameters $\alpha$’s, and $o_{i,t}$ indicates the proportion of time for which a camera trap is deployed at site $i$ in year $t$. Note that we assumed to have multiple camera traps deployed at each site to allow replicates of the observations (pollock1982capture); these camera traps need to function independently, meaning that capturing the animal by one camera trap should not influence the function of other camera traps. Also note that $r_{i,t,j}$ represents average number of detections, rather than a probability, when an activate individual spent unit time at a site that had unit survey effort $o_{i,t}$, and thus is not necessarily upper bounded by 1. Also, because we modelled all possible sites, $o_{i,t}$ allows the situation that camera traps are not necessarily deployed at every site and they can be deployed for different amount of time.
2.2 Simulation design
During the simulations, we simulated a total of 200 individuals, including individuals that actually existed and individuals that did not exist, meaning that we used data augmentation (royle2007analysis; lewis2023climate). More specifically, the parameter $\psi$ governs the number of individuals that actually exist in the first year of the study. In later years, some non-existing individuals can be recruited into the population, the probability of which is governed by $\gamma$. Eventually, there will be some individuals that never exist during the entire study period, and these individuals are ”augmented”. Being able to estimate the above-mentioned parameters indicates that data augmentation does not induce systematic bias when a proper number is selected for the augmented population size, shown by previous studies (royle2007analysis; milleret2023estimating) and the current study (see results). In practice, we recommend selecting a number that is a few times of the potential population size. Selecting a small number, which will be indicated by an estimated $\psi$ that is close to 1, can lead to underestimation of population size. On the other hand, selecting an extremely large number can make the computation inefficient and lead to small $\psi$ and $\gamma$ values that are difficult to estimate.
We simulated a landscape with 500 sites within a 20-unit wide and 16-unit long area during a 6-year period of time; the landscape was characterized by two spatially and temporally varying covariates, both of which were randomly generated from a standard Normal distribution. We assumed that territory centroids were uniformly distributed in the landscape in the first year. We set $\kappa$ to 2.5 to allow a moderate size of territory. We set $\beta^{[\upsilon]}_{1}$ to 0.6 and $\beta^{[\upsilon]}_{2}$ to -0.3 to allow opposite effects of the two covariates on habitat use. We set $\psi$ to 0.4, meaning that more than half of the individuals did not exist in the first year of the study. We set $\beta^{[\phi]}_{0}$ to 1 and $\beta^{[\gamma]}_{0}$ to -1 to allow a relatively high apparent survival and relatively low recruitment. We set $\beta^{[\phi]}_{density}$ to -0.4 and $\beta^{[\gamma]}_{density}$ to -0.6 to allow negative density dependence on both apparent survival and recruitment. We set $\sigma$ to 2.5 to allow a relatively short distance of dispersal movement. For the observation processes, we considered two observation covariates which again were randomly generated from a standard Normal distribution. We set $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$ to 0.7, -0.6, 0.3, respectively. For cameras in the middle part (16-unit wide and 12-unit long) of the study area, $o_{i,t}$ was randomly generated from a uniform distribution between 0.8 and 1; $o_{i,t}$ was set to 0 to represent the situation that cameras were not deployed outside the middle part of the study area.
[…]
2.3 Model evaluation
We wrote our own MCMC algorithm in R (r2024r). We repeated the simulations for 20 times with newly generated data each time. We ran 50,000 iterations, including 25,000 burn-in, in 3 chains for each simulation. We pooled all posterior samples from all simulations to make violin plots, which also provided key statistics including median, 50% Credible Interval (CI) and 95% CI. We considered a parameter estimate to be biased if the 50% CI does not cover the true value of the parameter used in the simulation.
2.4 Real data analysis
We applied our proposed method on nine years tiger survey data from India by karanth2006assessing. The published version of the dataset by tigerdata contained ten primary periods of surveys, however the first two surveys were conducted within the same year. We therefore removed the data from the second survey to constrain the midpoints of all surveys in the first half of the years for consistency. We delineated the study area into 342 grid cells that are 1 $km^{2}$ in size. When there were multiple (maximally 7) cameras deployed at the same sites, we considered them as spatial replicates. There were in total 74 individuals captured during these nine surveys, and we augmented our data to have 200 potential individuals. Survey effort, measured by days in each primary period a camera was active, ranged from 1 to 46 days. We divided these active days by the maximum active day (i.e., 46) to calculate the survey effort $o_{i,t}$, while using 46 as the time period covered by the survey $U$. The published version of the dataset does not contain any environmental variables, therefore we did not include any environmental covariates in our analysis. We ran three chains with 50,000 iterations including 10,000 burn-in in each chain. The convergence of the MCMC computing was checked using the diagnostic plot of the posterior samples and Brooks–Gelman–Rubin Diagnostics.
3 Results
3.1 Simulation
The simulations showed that both the simpler model that only allows density dependence on recruitment but not apparent survival and does not allow dispersal movement between years and the more complex model that allows density dependence on both recruitment and apparent survival and dispersal movement between years can recover parameters related to habitat use ($\kappa$, $\beta^{[\upsilon]}$’s), initial existence ($\psi$) and detection ($\alpha$’s) without evident bias (Figure 2). However, both models underestimated the effects of density dependence on apparent survival ($\beta_{density}^{[\phi]}$) and recruitment ($\beta_{density}^{[\gamma]}$). The simpler model provided unbiased estimates for the mean of apparent survival ($\beta_{0}^{[\phi]}$) and recruitment ($\beta_{0}^{[\gamma]}$) at a logit scale, but the more complex model tended o have slightly biased estimates for these two parameters. The models also underestimated the standard deviation of Gaussian dispersal movement ($\sigma$).
Figure 2: Violin plots showing the combined posterior samples from 20 simulations for the parameters of (M1) a SCR model that only allows density dependence on recruitment but not apparent survival and does not allow dispersal movement between years, and (M2) a SCR model that allows density dependence on both recruitment and apparent survival and dispersal movement.
3.2 Real data
Using our SCR model, we estimated parameters related to habitat use, density dependence apparent survival and recruitment, dispersal movement, and observation processes (Table.1). In particular, we found little to no effect of density dependence on apparent survival ($\beta^{[\phi]}_{density}$: median -0.01, 80% Credible Interval -0.13, 0.13), and a negative effect of density dependence on recruitment ($\beta^{[\gamma]}_{density}$: median -0.15, 80% Credible Interval -0.32, -0.01).
Parameter Median 80% CI
$\kappa$ 0.61 [0.53, 0.68]
$\psi$ 0.15 [0.09, 0.22]
$\beta^{[\phi]}_{0}$ 0.44 [-0.11, 1.08]
$\beta^{[\phi]}_{\text{density}} -$0.01 [-0.13, 0.13]
$\beta^{[\gamma]}_{0} -$1.04 [-2.02, -0.12]
$\beta^{[\gamma]}_{\text{density}} -$0.15 [-0.32, -0.01]
$\sigma$ 2.89 [2.49, 3.35]
pp 0.51 [0.44, 0.59]
Table 1: Posterior medians and 80% Credible Intervals of the parameters of the SCR model used in the tiger case study.
4 Discussion
In this study, we extended SCR models (e.g., royle2013integrating; royle2018unifying) to estimate the true amount of time of habitat use, an information that is often of interest in ecological research and conservation planning (hull2016habitat; dawson2013habitat). Using both simulations and a case study, we showed that it is possible to separate true habitat use from the observed number of captures under a hierarchical Bayesian Framework. A key component of our model is to allow habitat use be a function of both habitat heterogeneity and distance from territory centroid. Therefore, we highly recommend researchers to obtain and incorporate covariate information to maximize the utility of our model. Another key component of our model is to assume that the total amount of time covered by the survey, $U$, is known, under the assumption of closed population within each primary period. We recommend researchers to carefully select the primary period of their survey to meet this assumption. Further studies could also consider extending the model to allow the violation of this assumption due to, for example, death of an individual during the primary period.
Figure 3: The true (blue solid point: observed individual, blue broken point: unobserved individual) and estimated territory centroids (yellow cloud: observed individual, red cloud: unobserved individual) in one year of a spatial capture-recapture study. Grey crosses represent the locations of camera traps.
Despite the accurate information of habitat use provided by this SCR model, we found it still difficult to estimate the effect of density dependence on apparent survival and recruitment with this model. So far, density dependence is mostly considered at the population level in SCR models (e.g., chandler2014spatially); when density dependence is considered at individual level in SCR models, only the locations of territory centroids, but not the information of habitat use, were considered (milleret2023estimating). The shortcoming of doing so is that individuals will not influence each other through density dependent processes if their territory centroids do not fall at the same site, despite the fact that their territories can still overlap with each other to induce competition. Therefore, we attempted to develop a model that links habitat use information with density dependence. The difficulties in estimating the effect of density dependence in this model is likely due to the fact that the current SCR model can only identify the locations of territory centroids of observed individuals, but not unobserved individuals without linking the locations of territory centroids to environmental covariates (Figure 3). Because the locations of unobserved individuals are misidentified, the model mis-specifies local density, and thus underestimates the effect of density dependence on vital rates. It is likely for the same reason that the model also provides biased estimates of the movement parameter. In the case study, our model identifies no density dependence on apparent survival and a negative density dependence on recruitment. Because of the model’s tendency to underestimate density dependence, we cannot conclude that there is no density dependence on apparent survival, and studies have suggested that density dependent apparent survival is possible for tigers (bisht2019demography). On the other hand, the negative density dependence on recruitment seems to actually exit (although its strength may be underestimated), which indicates that the tigers may have a lower reproduction and/or immigration rate in areas used by more individuals. For instance, tigers are reported to be reluctant to move into a crowded area to prevent inbreeding depression (smith1991contribution).
Overall, our results are in line with the findings of previous studies that showed the limits of the current generation of SCR models (sun2024importance), and it remains a challenge to link habitat use with vital rates at individual level in SCR models. Therefore, we suggest researcher taking cautions when using this and other SCR models to evaluate density dependence. Further development of SCR models is needed to better identify the locations of individuals, probably through linking them to environmental covariates (e.g., borchers2008spatially) or the integration of other data resources with SCR data (e.g., sun2019incorporating).
Our study nonetheless showed the importance and possibility of separating true habitat use with observation processes in SCR models. Our modeling approach can be used in a broad range of species and populations as long as SCR surveys are applicable. Our model can also be easily adapted to a variety of sampling designs (dupont2021optimal; sun2014trap) or modified to include additional model components such as sex structures (augustine2020sex). Thus, our model is broadly transferable under the framework of spatial statistic trinity.